OpenStudy (anonymous):

The area of a regular hexagon is 40 cm^2. What is the area of of a regular hexagon with sides five times as long?

OpenStudy (anonymous):

I got 200 but I wanted to make sure I was correct.

OpenStudy (anonymous):

Uh..Okay.

OpenStudy (linyu):

so first find the side of a regular hexagon than use the formula again by letting the side = 5 times of a regular hexagon \[40 = x^23\sqrt{3}/2\] \[x = \sqrt{\frac{40}{ 3\sqrt{3}/2}}\] \[x = 4\sqrt{5}(3^{3/4})\] \[NewHexagonArea = 3\sqrt{3}/2*(4\sqrt{5}(3^{3/4}))^2\]

OpenStudy (linyu):

sorry it's \[Area = side^2 * 3\sqrt{3}/2\]

OpenStudy (anonymous):

Okay..I somewhat understand.

OpenStudy (linyu):

\[3\sqrt{3}/2*(5*4\sqrt{5}/3^{3/4})^2=1000\] i missed 5 on last solution.

OpenStudy (anonymous):

Thank you!

OpenStudy (linyu):

@Burningitdown you're very welcome